OpenStudy (anonymous):

PLEASE PLEASE PLEASE HELP! Complete the proof of the Quotient Rule: logb x/a = logb x – logb a. Let m = logb a and n = logb x. Then a = and x = By the Quotient Property of Exponents, x/a = Take logb of both sides to get logb x/a = Re-substitute for m and n to get logb x/a =

OpenStudy (ash2326):

do you know how log is defined? if you have \[\log_x y =z\] what's y here?

OpenStudy (ash2326):

@natasha.aries ???

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

@ash2326 sorry my computer was working

OpenStudy (anonymous):

the answer is probably over 9000

OpenStudy (anonymous):

its a proof though

OpenStudy (anonymous):

logb x/a = logb x – logb a. \(\large log_b \frac{x}{a} = log_bx-log_ba\) Let \(\large m=log_ba \) and \(\large n=log_bx \) so \(\large a=b^m \) and \(\large x=b^n \) so \(\large \frac{x}{a}=\frac{b^n}{b^m} =b^{n-m}\) therefore, \(\large log_b\frac{x}{a}=n-m=log_bx-log_ba \)

OpenStudy (anonymous):

thankss!